A 70kg woman dives from a 18m cliff. What is her velocity 5m into the dive (in m/s)? Assume no external forces are applied during diving.
Answer #1

Here,

let the velocity is v

using third equation of motion

v^2 - u^2 = 2 * a * d

v^2 - 0^2 = 2 * 9.8 * 5

v = 9.9 m/s

the velocity 5 m into the dive is 9.9 m/s
if we take distance = 13 m

then

v^2 - 0^2 = 2 * 9.8 * 13

v = 15.96 m/s

the speed is 15.96 m/s